OpenStudy (heidi710):

you toss 2 coins, Find P(at least 1 head is tossed) Please help!

OpenStudy (theweeknd16):

Tossing two coins, there are four possible outcomes: HH HT TH TT Three of the four have at least one head. That means your probability is 3/4, or 75%.

OpenStudy (heidi710):

thank you (:

OpenStudy (theweeknd16):

No problemo.

OpenStudy (trojanpoem):

at least 1 head is tossed is the same as not 2 tails) P(TT)^c = 1- P(TT) = 1- 0.5 * 0.5 = 3/4 This would be useful if the question were. if a coin was tossed 100000 times get the probability of at least one head. no way you're counting that.